OpenStudy (anonymous):

solve the equation for t 3t - 6t + 8 = 7t

OpenStudy (anonymous):

t = 8/10 or 4/5

OpenStudy (anonymous):

if by t 3t you mean 3t^2 then you need to use the quadratic formula. If you need to use that formula then you'll get 3.703 and .6301

OpenStudy (anonymous):

3t-6t-7t=-8 3t-13t=-8 -10t=-8 10t=8 t=8/10 t=4/5